Q.20 An enzyme catalyzed reaction (following Michaelis-Menten kinetics) exhibits maximum reaction
velocity (Vm) of 75 nmol l
-1
min-1
.

The enzyme at a substrate concentration of 1.0x10-4
M shows the
initial reaction velocity of 60 nmol l
-1
min-1
. The Km value of the enzyme in molar concentration
(M) is
(A) 2.5 x 10-5
(B) 5.0 x 10-5

(C) 2.5 x 10-4
(D) 5.0 x 10-4
1
Expert's answer
2014-01-13T05:14:28-0500
V= Vm*S/(S+Km)
V*S+V*Km=Vm*S
Km=S*(Vm-V)/ V
Km=2.5*10^-5 M
Variant A is correct